In a chemistry class, 4 liters of a 4% silver iodine solution must be mixed with a 10% solution to get a 6% solution. How many liters of the 10% solution are needed?
Answer by stanbon(75887)

In a chemistry class, 4 liters of a 4% silver iodine solution must be mixed with a 10% solution to get a 6% solution. How many liters of the 10% solution are needed?
------
Equation:
active + active = active
0.04*4 + 0.10x = 0.06(4+x)
----
Multiply thru by 100 to get:
4*4 + 10x = 6*4 + 6x
---
4x = 2*4
x = 2 liters (amt. of 10% solution needed)
===============
Cheers,
Stan H.